JET-COOLED CAVITY RING-DOWN SPECTROSCOPY OF THE Ã^{2}E′′-X̃^{2}A′_{2} VIBRONIC TRANSITION OF NO_{3}

\underline{MING-WEI CHEN}, GABRIEL M. P. JUST^{a}, TERRANCE CODD, and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210.

The three energetically lowest electronic states (X̃ ^{2}A′_{2}, Ã ^{2}E′′, and B̃ ^{2}E′) of NO_{3} are strongly coupled by vibronic interactions and have been treated in considerable detail theoretically.^{b} Corresponding experimental characterization of the interaction is much less detailed. Previous experimental results primarily consist of IR measurements of vibrational transitions in the ground state.^{c}^{d} In addition, the electronically forbidden Ã-X̃ transition has been observed in ambient temperature CRDS studies.^{e}^{f} A slit-jet nozzle with a high voltage pulsed discharge has been applied to produce the NO_{3} radical by dissociating the N-O bond of N_{2}O_{5}, and the jet-cooled NO_{3} CRDS absorption spectrum has been successfully observed with a high-resolution laser source (Δν ≈250MHz, intrinsic resolution considering the instrumental linewidth and the residual Doppler broadening in the jet). The 4_{0}^{1} band (parallel band) shows complex rotational structure which is presently being analyzed. The 2_{0}^{1} band has also been measured as an example of a perpendicular band. Besides the ν_{2} and ν_{4} vibronic bands, the vibronically forbidden origin band (0_{0}^{0} band) has been recorded under the same experimental conditions. The weakly observed Ã-X̃ origin band is likely either a magnetic dipole or an electric quardrupole transition.

apresent address: Lawrence Berkeley National Laboratory, Lawrence Berkeley National Laboratory

bJ. F. Stanton, J. Chem. Phys., \underline{126}, 134309 (2007)

cK. Kawaguchi, E. Hirota, T. Ishiwata, and I. Tanaka, J. Chem. Phys., \underline{93}, 951 (1990)

dK. Kawaguchi, T. Ishiwata, E. Hirota, and I. Tanaka, Chem. Phys., \underline{231}, 193 (1998)

eA. Deev, J. Sommar, and M. Okumura, J. Chem. Phys., \underline{122}, 224305 (2005)

fE. Hirota, T. Ishiwata, K. Kawaguchi, M. Fujitake, N. Ohashi, I. Tanaka, J. Chem. Phys., \underline{107}, 2829 (1997)